# Terminal Velocity Matching

### v2

@@ 6 Related Works
Diffusion and Flow Matching. Diffusion models (sohl2015deep; ho2020denoising; song2020score) learn generative models by reversing stochastic processes, while Flow Matching (liu2022flow; lipman2022flow) generalizes this to arbitrary priors with simplified training. Both approaches ultimately solve ODEs with neural networks during sampling.
One-Step and Few-Step Models from Scratch. To address slow inference from ODE simulation, recent methods aim for few-step generation in a single training stage. Consistency models (song2023consistency; lu2024simplifying) parameterize networks to represent ODE integrals but cannot jump between arbitrary timesteps without injecting additional noise, which can limit multi-step performance. Two-time conditioned approaches enable arbitrary timestep transitions: IMM (zhou2025inductive) provides distribution consistency via Maximum Mean Discrepancy but requires multiple particles; MeanFlow (geng2025mean) and Flow Map Matching (boffi2024flow) match trajectory derivatives but lack distributional guarantees. Other variants bypass differentiation via Monte Carlo (liu2025learning) or combine distillation with FM (frans2024one; boffi2025build).
Unlike these methods, TVM regularizes path behavior at the terminal time rather than initial time and provides explicit $2$-Wasserstein bounds. While sCT and MeanFlow only compute forward JVP, TVM uniquely supports backward passes through the JVP computation, enabling full gradient flow for the terminal velocity objective. These innovations drive both our theoretical insights and architectural improvements.
@@ 7.1 Image Generation
ImageNet-${\mkern-1.0mu\times\mkern-1.0mu}256$. We present quantitative results in Table 1 under FID (heusel2017gans). We adopt the default DiT-XL/2 architecture (peebles2023scalable) and inject tst-s as the second timestep, following IMM (zhou2025inductive) and MeanFlow (geng2025mean). We additionally employ our semi-Lipschitz control techniques for training stability or we notice activation explosion as described in Figure 4, and we train with constantly sampled CFG, i.e. models with $w=2$ and $w=1.75$ are two different models trained from scratch. We describe additional training details in Appendix F. Our method achieves state-of-the-art 1-NFE FID among methods trained from scratch, outperforming MeanFlow and IMM. With CFG $w=2$, TVM can achieve noticeable improvements over MeanFlow, e,g, 3.29 FID vs. 3.43 FID for 1-NFE and similarly for 2-NFE. With 4 NFEs, $w=1.75$ also exceeds 500-NFE diffusion baselines. We additionally show qualitative 1-NFE samples on the right of Figure 6.
ImageNet-${\mkern-1.0mu\times\mkern-1.0mu}512$. We train with the same settings as in $256{\mkern-1.0mu\times\mkern-1.0mu}256$-resolution and we show the FID scores in Table 2. We rerun MeanFlow using the same settings as in ImageNet-$256{\mkern-1.0mu\times\mkern-1.0mu}256$ as our baseline in addition to sCT (lu2024simplifying) under similar model sizes. TVM again outperforms sCT and MeanFlow in 1-NFE and 2-NFE regime. Notably, TVM-XL/2 outperforms sCT-XL with 1.1B parameters, highlighting TVM’s more optimal use of model capacity in fitting the image distribution. Moreover, with $w=2.25$, TVM with 4-NFE can match 500-NFE DiT-XL/2 baseline in performance, further demonstrating the scalability of our algorithm to higher resolution.
@@ 7.2 Discussion on Training Advantages
Single sample objective. Unlike IMM (zhou2025inductive) which uses more than 4 samples to calculate its loss, we use a single sample to for loss calculation without losing a distribution-matching interpretation. This also allows the objective to be scaled to large models and high-dimensional datasets where batch size on each GPU is constrained to be 1.
Training with random CFG. Our construction allows us to randomly sample CFG scale during training without collapse. We attribute this stability to our JVP being only calculated w.r.t. $s$ which is invariant to starting position ${\mathbf{x}}_{t}$ and time $t$. In contrast, CT (song2023consistency; lu2024simplifying) and MeanFlow (geng2025mean) require velocity ${\mathbf{u}}({\mathbf{x}}_{t},t)$ to be used in the JVP calculation. In the case of random CFG, this velocity can vary widely in magnitude which, if propagated through JVP, can cause wide fluctuation in gradient norm (see left two in Figure 7) and causes training instability. Our method, in comparison, enjoys much smoother gradient norm and ${\mathbf{u}}({\mathbf{x}}_{t},t)$ norm, and successfully converges even in the presence of random CFG. We note that random sampling of CFG is not optimal as some CFG scales experience degradation in FID during training, and constant CFG performs better in comparison. We postulate that the under-performance of random CFG is due to limited capacity of the network and the $1/w^{2}$ factor that downweights high CFG. This phenoemenon is similarly observed in CFG-conditioned FM training (see Appendix F.3) and we leave any improved design to future work.
@@ 7.3 Ablation Studies
EMA target rate $\gamma$. The target EMA weight $\theta^{*}$ plays a significant role in accelerating convergence of the model. Shown in Table 5(b), non-EMA target, i.e. $\gamma=0$, noticeably lags behind $\gamma>0$ alternatives. However, too large of a $\gamma$, e.g. 0.9999, also causes instability because of the overly slow target update. A sweet spot exists around $\gamma=0.99$ which we use as default. Besides attribute its success to variance reduction because EMA’s slower weight update implies much lower optimization noise. In addition, EMA is commonly used to evaluate diffusion models for its quality boost (song2020score), so being the optimization target also provides better learning signal to the model.
@@
Probability for $t=s$. Inspired by MeanFlow (geng2025mean), we investigate whether setting $t=s$ (when it reduces to pure FM training) is helpful for overall performance. We find that $0\%$ actually degrades 1-NFE performance while it marginally improves 2-NFE performance. This tradoff persists throughout the training but we observe diminishing return as training goes on. Therefore, we do not find this practice helpful in general and leave it out of our design space in general.
@@ A.4 Main Theorem
where $\tilde{C}$ is some non-optimizable constant (lipman2022flow). This is also a classical result connecting score matching and denoising score matching (vincent2011connection).
@@ Appendix C Flash Attention JVP with Backward Pass
Computing the Jacobian-Vector Product (JVP) of SDPA is even more demanding, typically requiring about three times the cost of the standard forward pass. Flash attention (dao2022flashattention) fuses the matrix multiplication with an online softmax operation (milakov2018onlinesoftmax), thereby eliminating the need to store the intermediate $QK^{\top}$ matrix in GPU memory. Subsequent work has shown that JVP SDPA can also be implemented in a FlashAttention-style manner, where both, primal SDPA and JVP SDPA are computed jointly to avoid redundant computation (lu2024simplifying).
Building on these ideas, we implement efficient JVP SDPA forward and backward kernels in Triton. We first take inspiration from open-source implementations without backward support [^4] https://github.com/Ryu1845/min-sCM/blob/main/standalone_multihead_jvp_test.py. And the additional backward pass through the standard (“primal”) SDPA is handled independently using the open-source implementation from (dao2022flashattention). To obtain full gradients with respect to $Q$, $K$, and $V$, we combine the input gradients from both backward passes.
@@ Appendix D Training Algorithm
Appendix E Relation to Prior Works
@@ F.1 Architecture and Optimization
VAE. We follow zhou2025inductive for the VAE setting, which uses the standard Stable Diffusion VAE (rombach2022high) but with a different scale and shift. Please refer to the paper for details.
Architecture. All architecture decisions follow DiT (peebles2023scalable) except for the changes described in the main text. For our XL-sized model, we follow DiT-XL and use 1152 hidden size but use 18 heads instead of 16 heads. This is purely for efficiency reasons because 18 heads under 1152 total hidden size implies head dimension is 64, while the original 16 heads result in head dimension 72. Flash attention JVP’s runtime is sensitive to redundancy in memory allocations. As 64 is a power of 2 our kernel can fully allocate appropriately sized CUDA blocks, while 72 leaves significant chunks unused. We observe that the original 16-head decision is $\times 1.25$ slower than the 18-head variant. In comparing FID of the two versions, we observe they perform similarly throughout training.
Following zhou2025inductive, we use tst-s as our second time condition into the architecture rather than directly injecting $s$. For injecting $w$, we follow chen2025visual and use $\beta=1/w$ as our condition, and if random CFG is used training, we sample $\beta\sim{\mathcal{U}}(\frac{1}{w_{\text{max}}},\frac{1}{w_{\text{min}}})$ and set $w=1/\beta$. Note that chen2025visual uses $\beta\sim{\mathcal{U}}(0,1)$ which amounts to $w_{\text{min}}=1$ and $w_{\text{max}}=\infty$, but arbitrarily large $w$ is never used in practice so $w_{\text{max}}$ can be set to a realistic finite value.
@@ F.2 Details on Random CFG with MeanFlow
In MeanFlow (geng2025mean), the authors introduce a mixing scale $\kappa$ such that the field with guidance scale $w$ is given by
@@ F.3 CFG-Conditioned Flow Matching
As in our method, we similarly observe tradeoff in FID if FM is trained to condition on CFG scale $w$ with randomly sampled $w$ during training (chen2025visual). During inference time, $w$ is injected into the network so that the CFG velocity field can be approximated by a single forward call. We inject $w$ using positional embedding just like the diffusion time, and during training we sample $\beta\sim{\mathcal{U}}(0,1)$ and set $w=1/\beta$, following chen2025visual. We show in Figure 11 that as the model trains, the FID of $w=1.5$ decreases but $w=2$ increases for later training steps. This tradeoff is similarly observed in our method as presented in the main text.

### v3

@@ 6 Related Works
Diffusion and Flow Matching. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020) learn generative models by reversing stochastic processes, while Flow Matching (Liu et al., 2022; Lipman et al., 2022) generalizes this to arbitrary priors with simplified training. Both approaches ultimately solve ODEs with neural networks during sampling.
One-Step and Few-Step Models from Scratch. To address slow inference from ODE simulation, recent methods aim for few-step generation in a single training stage. Consistency models (Song et al., 2023; Lu and Song, 2024) parameterize networks to represent ODE integrals but cannot jump between arbitrary timesteps without injecting additional noise, which can limit multi-step performance. Two-time conditioned approaches enable arbitrary timestep transitions: IMM (Zhou et al., 2025) provides distribution consistency via Maximum Mean Discrepancy but requires multiple particles; MeanFlow (Geng et al., 2025) and Flow Map Matching (Boffi et al., 2024) match trajectory derivatives but lack distributional guarantees. Other variants bypass differentiation via Monte Carlo (Liu and Yue, 2025) or combine distillation with FM (Frans et al., 2024).
Unlike these methods, TVM regularizes path behavior at the terminal time rather than initial time and provides explicit $2$-Wasserstein bounds. While sCT and MeanFlow only compute forward JVP, TVM uniquely supports backward passes through the JVP computation, These innovations drive both our theoretical insights and architectural improvements. Another recent work (Boffi et al., 2025) has proposed a unifying perspective on training flow maps from scratch; it encompasses many approaches such as Shortcut (Frans et al., 2024), CM (Song et al., 2023; Lu and Song, 2024), MeanFlow (Geng et al., 2025) and shares similar theoretical insights as TVM.
@@ 7.1 Image Generation
ImageNet-${\mkern-1.0mu\times\mkern-1.0mu}256$. We present quantitative results in Table 1 under FID (Heusel et al., 2017). We adopt the default DiT-XL/2 architecture (Peebles and Xie, 2023) and inject tst-s as the second timestep, following IMM (Zhou et al., 2025) and MeanFlow (Geng et al., 2025). We additionally employ our semi-Lipschitz control techniques for training stability or we notice activation explosion as described in Figure 4, and we train with constantly sampled CFG, i.e. models with $w=2$ and $w=1.75$ are two different models trained from scratch. We describe additional training details in Appendix F. Our method achieves state-of-the-art 1-NFE FID among methods trained from scratch, outperforming MeanFlow and IMM. With CFG $w=2$, TVM can achieve noticeable improvements over MeanFlow, e,g, 3.29 FID vs. 3.43 FID for 1-NFE and similarly for 2-NFE. With 4 NFEs, $w=1.75$ also exceeds 500-NFE diffusion baselines. We additionally show qualitative 1-NFE samples on the right of Figure 6.
ImageNet-${\mkern-1.0mu\times\mkern-1.0mu}512$. We train with the same settings as in $256{\mkern-1.0mu\times\mkern-1.0mu}256$-resolution and we show the FID scores in Table 2. We rerun MeanFlow using the same settings as in ImageNet-$256{\mkern-1.0mu\times\mkern-1.0mu}256$ as our baseline in addition to sCT (Lu and Song, 2024) under similar model sizes. TVM again outperforms sCT and MeanFlow in 1-NFE and 2-NFE regime. Notably, TVM-XL/2 outperforms sCT-XL with 1.1B parameters, highlighting TVM’s more optimal use of model capacity in fitting the image distribution. Moreover, with $w=2.25$, TVM with 4-NFE can match 500-NFE DiT-XL/2 baseline in performance, further demonstrating the scalability of our algorithm to higher resolution.
@@ 7.2 Discussion on Training Advantages
Single sample objective. Unlike IMM (Zhou et al., 2025) which uses more than 4 samples to calculate its loss, we use a single sample to for loss calculation without losing a distribution-matching interpretation. This also allows the objective to be scaled to large models and high-dimensional datasets where batch size on each GPU is constrained to be 1.
Training with random CFG. Our construction allows us to randomly sample CFG scale during training without collapse. We attribute this stability to our JVP being only calculated w.r.t. $s$ which is invariant to starting position ${\mathbf{x}}_{t}$ and time $t$. In contrast, CT (Song et al., 2023; Lu and Song, 2024) and MeanFlow (Geng et al., 2025) require velocity ${\mathbf{u}}({\mathbf{x}}_{t},t)$ to be used in the JVP calculation. In the case of random CFG, this velocity can vary widely in magnitude which, if propagated through JVP, can cause wide fluctuation in gradient norm (see left two in Figure 7) and causes training instability. Our method, in comparison, enjoys much smoother gradient norm and ${\mathbf{u}}({\mathbf{x}}_{t},t)$ norm, and successfully converges even in the presence of random CFG. We note that random sampling of CFG is not optimal as some CFG scales experience degradation in FID during training, and constant CFG performs better in comparison. We postulate that the under-performance of random CFG is due to limited capacity of the network and the $1/w^{2}$ factor that downweights high CFG. This phenoemenon is similarly observed in CFG-conditioned FM training (see Appendix F.3) and we leave any improved design to future work.
@@ 7.3 Ablation Studies
EMA target rate $\gamma$. The target EMA weight $\theta^{*}$ plays a significant role in accelerating convergence of the model. Shown in Table 5(b), non-EMA target, i.e. $\gamma=0$, noticeably lags behind $\gamma>0$ alternatives. However, too large of a $\gamma$, e.g. 0.9999, also causes instability because of the overly slow target update. A sweet spot exists around $\gamma=0.99$ which we use as default. Besides attribute its success to variance reduction because EMA’s slower weight update implies much lower optimization noise. In addition, EMA is commonly used to evaluate diffusion models for its quality boost (Song et al., 2020), so being the optimization target also provides better learning signal to the model.
@@
Probability for $t=s$. Inspired by MeanFlow (Geng et al., 2025), we investigate whether setting $t=s$ (when it reduces to pure FM training) is helpful for overall performance. We find that $0\%$ actually degrades 1-NFE performance while it marginally improves 2-NFE performance. This tradoff persists throughout the training but we observe diminishing return as training goes on. Therefore, we do not find this practice helpful in general and leave it out of our design space in general.
@@ A.4 Main Theorem
where $\tilde{C}$ is some non-optimizable constant (Lipman et al., 2022). This is also a classical result connecting score matching and denoising score matching (Vincent, 2011).
@@ Appendix C Flash Attention JVP with Backward Pass
Computing the Jacobian-Vector Product (JVP) of SDPA is even more demanding, typically requiring about three times the cost of the standard forward pass. Flash attention (Dao et al., 2022) fuses the matrix multiplication with an online softmax operation (Milakov and Gimelshein, 2018), thereby eliminating the need to store the intermediate $QK^{\top}$ matrix in GPU memory. Subsequent work has shown that JVP SDPA can also be implemented in a FlashAttention-style manner, where both, primal SDPA and JVP SDPA are computed jointly to avoid redundant computation (Lu and Song, 2024).
Building on these ideas, we implement efficient JVP SDPA forward and backward kernels in Triton. We first take inspiration from open-source implementations without backward support [^4] https://github.com/Ryu1845/min-sCM/blob/main/standalone_multihead_jvp_test.py. And the additional backward pass through the standard (“primal”) SDPA is handled independently using the open-source implementation from (Dao et al., 2022). To obtain full gradients with respect to $Q$, $K$, and $V$, we combine the input gradients from both backward passes.
@@ Appendix D Training Algorithm
Appendix E Relation to Other Works
@@ E.1 MeanFlow
E.2 Flow Map Self-Distillation
Boffi et al. (2025) proposes three different self-distillation approaches for training flow maps from scratch.
@@ F.1 Architecture and Optimization
VAE. We follow Zhou et al. (2025) for the VAE setting, which uses the standard Stable Diffusion VAE (Rombach et al., 2022) but with a different scale and shift. Please refer to the paper for details.
Architecture. All architecture decisions follow DiT (Peebles and Xie, 2023) except for the changes described in the main text. For our XL-sized model, we follow DiT-XL and use 1152 hidden size but use 18 heads instead of 16 heads. This is purely for efficiency reasons because 18 heads under 1152 total hidden size implies head dimension is 64, while the original 16 heads result in head dimension 72. Flash attention JVP’s runtime is sensitive to redundancy in memory allocations. As 64 is a power of 2 our kernel can fully allocate appropriately sized CUDA blocks, while 72 leaves significant chunks unused. We observe that the original 16-head decision is $\times 1.25$ slower than the 18-head variant. In comparing FID of the two versions, we observe they perform similarly throughout training.
Following Zhou et al. (2025), we use tst-s as our second time condition into the architecture rather than directly injecting $s$. For injecting $w$, we follow Chen et al. (2025) and use $\beta=1/w$ as our condition, and if random CFG is used training, we sample $\beta\sim{\mathcal{U}}(\frac{1}{w_{\text{max}}},\frac{1}{w_{\text{min}}})$ and set $w=1/\beta$. Note that Chen et al. (2025) uses $\beta\sim{\mathcal{U}}(0,1)$ which amounts to $w_{\text{min}}=1$ and $w_{\text{max}}=\infty$, but arbitrarily large $w$ is never used in practice so $w_{\text{max}}$ can be set to a realistic finite value.
@@ F.2 Details on Random CFG with MeanFlow
In MeanFlow (Geng et al., 2025), the authors introduce a mixing scale $\kappa$ such that the field with guidance scale $w$ is given by
@@ F.3 CFG-Conditioned Flow Matching
As in our method, we similarly observe tradeoff in FID if FM is trained to condition on CFG scale $w$ with randomly sampled $w$ during training (Chen et al., 2025). During inference time, $w$ is injected into the network so that the CFG velocity field can be approximated by a single forward call. We inject $w$ using positional embedding just like the diffusion time, and during training we sample $\beta\sim{\mathcal{U}}(0,1)$ and set $w=1/\beta$, following Chen et al. (2025). We show in Figure 11 that as the model trains, the FID of $w=1.5$ decreases but $w=2$ increases for later training steps. This tradeoff is similarly observed in our method as presented in the main text.